OpenStudy (anonymous):

Please help with these Algebra 2 questions. I'll give a medal. 1. What value completes the square for the expression? x^2+14x 2. Use the quadratic formula to solve the equation. x^2-7x-6=0 3. What are the solutions? 1/2x^2+2x+3=0

OpenStudy (shinalcantara):

Completing the square - This only applies if the coefficient of x^2 is 1. okay? 1. Identify the middle term-the term with x as the variable. (In this case, it's 14) 2. Divide the numerical coefficient of the middle term by 2 (14/2 = 7) 3. To get the third term, square the quotient from step 2. (7^2 = 49) Then you'll have it as \[x^2 + 14x +49\]

OpenStudy (shinalcantara):

Standard equation: \[ax^2 + bx + c = 0\] Quadratic Formula: \[x = \frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (shinalcantara):

\[x^2-7x-6=0\] \[a=1\] \[b=-7\] \[c=-6\]

OpenStudy (shinalcantara):

Just plug in the values to the formula

OpenStudy (shinalcantara):

is this\[\frac{1}{2}x^2+2x+3=0 \]

OpenStudy (shinalcantara):

@KatieBear ?

OpenStudy (anonymous):

yeah, it is. Sorry, I was eating breakfast. lol

OpenStudy (shinalcantara):

\[\frac{ 1 }{ 2 }x^2 + 2x + 3 = 0\] Multiply the whole equation by 2 \[x^2 + 4x + 6 = 0\] Use quadratic equation a = 1 b = 4 c = 6 --------------------- I'm pretty sure that the roots are imaginary xD

OpenStudy (shinalcantara):

*quadratic formula

OpenStudy (anonymous):

Thank you so much! I have a test today and now I understand the material so much better. :)

OpenStudy (shinalcantara):

yw =)